OpenStudy (anonymous):

if f(x)=ax+b where a and b are constants, find a and b for f(1)=-4 and f(3)=2

TheSmartOne (thesmartone):

So when x=1 y=-4 and when x=3 y=2

TheSmartOne (thesmartone):

f(1)=a(1)+b -4=a+b and f(3)=a(3)+b 2=3a+b

TheSmartOne (thesmartone):

So you need to solve the two equations by substitution or elimination -4=a+b 2=3a+b

TheSmartOne (thesmartone):

you can just subtract both of them and eliminate b and you will get -6=-2a

TheSmartOne (thesmartone):

divide by -2 and a=3

TheSmartOne (thesmartone):

now plug in a=3 and -4=3+b b=-7

TheSmartOne (thesmartone):

So your solutions are a=3 and b=-7 @philosophia

OpenStudy (anonymous):

thank youuu :) @TheSmartOne